Jacob wants to compute 23x23, he says that 20x20=400 and 3x3=9 so the product must be 400+-=409. What would you discuss with Jacob to reinforce that he is partly correct and show him what he has missed
Answer by Earlsdon(6294)

I would tell Jacob that 23 X 23 can be written as (20+3)X(20+3) and that he can use the FOIL method to perform this multiplication thus:
(23x23) = (20+3)(20+3) = (20X20)+(20X3)+(3X20)+(3x3) = 400+60+60+9 = 529